OpenStudy (anonymous):

area of surface obtained by revovling the graph y=x^4/4+1/8x^2 from 1 to 2?

OpenStudy (anonymous):

revolving around the x-axis or the y-axis?

OpenStudy (anonymous):

y axis

OpenStudy (anonymous):

m sorry

OpenStudy (anonymous):

Use the shell method, the volume of a shell \[ 2\pi x\left ( \frac 9 2 -y \right)dx \] 9/2 is for x =1

OpenStudy (anonymous):

\[2\pi x\left ( \frac 9 2 -\frac{x^4}{4}-\frac{x^2}{8} \right)dx=\\ 2\pi \left ( \frac 9 2 x -\frac{x^5}{4}-\frac{x^3}{8} \right)dx \] Integrate this quantity from x=1 to x=2

OpenStudy (anonymous):

9/2 is for x=2 (correction to a previous post)

OpenStudy (anonymous):

Your final answer should be \[\frac{117 \pi }{16} \]